what is: 1/2x +4 = 2/3
show steps

Answer by checkley71(8403)

1/2X+4=2/3 SUBTRACT 4 FROM BOTH SIDES.
X/2=2/3-4 COMBINE THE FRACTION & THE DIGIT.
X/2=(2-12)/3 REDUCE THE NUMERATOR.
X/2=-10/3 CROSS MULTIPLY.
3X=-20 DIVIDE BOTH SIDES BY 3.
X=-20/3 ANSWER.